OpenStudy (anonymous):

Choose the point that lies on the graph of the piecewise function. f of x equals negative the quantity of x plus 5 end quantity squared if x is less than or equal to negative 3 and equals 2x minus 5 if x is greater than negative 3.

OpenStudy (anonymous):

@neer2890

OpenStudy (neer2890):

write your question in mathematical form first.

OpenStudy (anonymous):

@neer2890

OpenStudy (anonymous):

(-2, 9) (-3, -11) (-7, -4) (-4, -13)

OpenStudy (neer2890):

you have any options or i have to choose it myself..?

OpenStudy (anonymous):

these r the points

OpenStudy (neer2890):

so we know that f(x)=y when the value of x is less than or equal to -3,we have to choose \[y=-(x+5)^{2}\] and when it is greater than -3 we have to choose 2x-5

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

Additionally, we know that the point which we are going to find will lie on graph, so the point should satisfy the equation of graph.

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

for e.g., (-2,9) here, x=-2 and y=9 so when x=-2, we have y=2x-5 should satisfy the equation. just put x=-2 y=2(-2)-5=-4-5=-9 but we have 9 as y coordinate so (-2,9) does not satisfy the equation. similarly, chk for other options

OpenStudy (anonymous):

ok thnx

OpenStudy (neer2890):

you're welcome...:) just check that you should get c as correct option when you solve this.

OpenStudy (anonymous):

ok and i think i got b

OpenStudy (neer2890):

no. you should get c. how did you get b..?

OpenStudy (anonymous):

2(-3)-5=-6-5=-11 and we have -11 as y coordinate

OpenStudy (neer2890):

ok. so when x=-3 you donot have to choose y= 2x-5 because it's not defined for x=-3 for x=-3, you can only use \[y=-(x+5)^{2}\]

OpenStudy (anonymous):

oh i got it

OpenStudy (neer2890):

ok... all the best

OpenStudy (anonymous):

thnx